# Pau S. Pujolas

### v1

@@ 4 Quantitative Model
where ${t_{ij}^{s}}$ is an iceberg trade cost, for example due to non-tariff trade barriers or freight costs, and $z_{j}^{s}$ is productivity. This means labor is given by
@@
Intermediate inputs used in each sector sum up to the total supply of intermediate inputs
$\sum_{s=1}^{S}\sum_{i=1}^{J}\left(x_{ij}^{sk}\right)=x_{j}^{k},$
and intermediate inputs plus consumption equal the total supply of the good
$x_{j}^{k}+c_{j}^{k}=y_{j}^{k}.$
@@ 5.1 Data on Flows and Tariffs
We use the 2016 Revision of the World Input-Output Database (WIOD), as described by Timmer et al. (2015) and Timmer et al. (2016), for data on industry-level trade flows and input-output flows. This dataset includes origin-destination-sector expenditures, $p_{ij}^{s}y_{ij}^{s}/\tau_{ij}^{s}$, net of tariffs, and sector-to-sector input-output flows, $p_{j}^{k}x_{j}^{sk}$. Note that in our model, $p_{ij}^{s}y_{ij}^{s}$  is expenditures from the consumer side, which include tariffs. The flows in the WIOD data do not include tariffs, therefore we need to add them in to get tariff-inclusive expenditures. We use data on tariff rates from the United Nations’ Trade Analysis Information System (TRAINS) database. We use effectively applied weighted average tariff rates as our measure for baseline tariffs, $\tau_{ij}^{s}$. Table 1 reports pre-trade war tariff rates for each of our tradable goods-producing industries. [^15] Construction is included the Goods-Producing Industries super-sector classification by the United States’ Bureau of Labor Statistics. We exclude construction from our grouping of tradable goods-producing industries, and include it as a service industry. Service industries are still tradable, but they are not subject to tariffs. The column "USA" reports tariff rates enacted by the United States for imports from China, and the column "CHN" reports tariff rates enacted by China for imports from the United States. [^16] The tariff data provided by TRAINS is available at the 2-digit International Standard Industrial Classification of All Economic Activities (ISIC) Rev 3 level, which we convert to match the ISIC Rev. 4 sectors in the WIOD using the concordance provided by Pahl et al. (2022).
@@ 5.2 Calibration
We take estimates of the sectoral elasticities, $\sigma_{s}$, from Fontagné et al. (2022), who provide sectoral elasticity estimates for each of the tradable WIOD sectors. Their estimates are based on product-level variation in tariff rates and trade-flows, which they pool by sector to estimate sectoral elasticities. They are unable to estimate elasticities for service sectors, since trade in service sectors is not subject to tariffs. We set elasticities for service industries equal to 5.5, which is near the median value of their product-level trade elasticities.
@@
We calibrate the aggregate trade balance directly from equation (21) and initial tariff revenues from equation (20). Using equation (12), we compute Income as
$\text{Income}_{i}=\sum_{s=1}^{S}p_{i}^{s}c_{i}^{s}=\sum_{s=1}^{S}\sum_{i=1}^{J}p_{ij}^{s}y_{ij}^{s}-\sum_{s=1}^{S}p_{i}^{s}x_{i}^{s},$ (23)
@@ 5.3 Computation of Optimal Tariffs
We focus our analysis on optimal tariff setting between two trading partners, $i$ and $j$, since the application motivating our analysis is a tariff war between two countries with a bilateral trade imbalance. To compute optimal tariffs, we solve the problem of a government choosing tariffs with a specific trading partner to maximize domestic welfare. Specifically, the government for country $i$ chooses tariffs in each sector, $\boldsymbol{\tau}_{i}\equiv\{\tau_{ij}^{s}\}_{s=1}^{S}$, to solve
@@
In general, there is no analytic solution for best response tariffs or Nash tariffs. Consequently, standard solution methods require evaluating counterfactual welfare for potential tariff candidates to find the candidates that numerically yield the highest welfare, which necessitates solving the model a large number of times. Past quantitative work has dealt with this computational intensity by focusing on small models, for example with one or two sectors (see He et al., 2017; or Chattopadhyay and Mitka, 2019), by focusing on environments where optimal tariff setting is either shown to be, or restricted to be, uniform across sectors (Caliendo et al., 2023); or by characterizing properties of optimal tariffs without calculating specific values (Costinot et al., 2020). The general approach to the problem of numerically solving for optimal tariff rates is given by Algorithm 1 (A more detailed approach is described by Algorithm A.1 in Appendix A). Running Algorithm 1 a single time with baseline tariffs yields optimal unilateral tariffs.
@@
Huang et al. (2013) show that, for certain parameterizations, it is possible to have multiple Nash equilibria in a simple two country two good CES framework. Particularly with low trade elasticities, they demonstrate the existence of Nash equilibria with extremely high tariff rates, including rates greater than 100,000 percent. For other parameterizations, the Nash equilibrium will be unique. To limit any potential impact of multiple equilibria, we restrict our analysis to tariff rates capped above at 400 percent. This rate exceeds the 99th percentile of industry level tariff rates in the TRAINS data, which is approximately 55 percent. This cap is also higher than the maximum reported ad valorem tariff rate of 350 percent set on an 8 digit Harmonized System product code by the United States in both 2014 and 2023, according to the United States International Trade Commission’s Tariff Database. [^19] https://dataweb.usitc.gov/tariff/annual We confirm that moderately lowering or raising this upper bound does not affect our quantitative results.
@@
We discuss our implementation of a hybrid genetic algorithm in Appendix A. As noted previously, there is nothing inherently unique about our chosen approach, and feasible alternatives exist, including random search, simulated annealing (Wu and Wang, 1998), particle swarm optimization (Kennedy and Eberhart, 1995), and hyperband optimization (Li et al., 2018). It should be noted that grid search is implausible due to the computational unfeasibility of evaluating all points on the grid, even for small grids. [^21] As noted in the introduction, even a discretized solution space of only 10 possible tariff rates for each sector yields 10 Sextillion possible combinations. We do not provide a detailed comparison on the relative performances of various global optimization heuristics, in part because each of them can be modified and adapted to perform better on specific problems. For instance, the genetic algorithm described in Appendix A will outperform a generic implementation of a genetic algorithm in terms of both accuracy and speed when applied to our problem of computing optimal tariff rates.
After convergence, we verify through exhaustive search that there are no single-sector deviations from Nash equilibrium tariffs that improve welfare. We evaluate counterfactual welfare changes for all tariffs between 0 and 25.00 percent to two decimal places for each sector. This encompasses the range of optimal tariffs we find, and we verify separately with a coarser grid that deviations with higher tariff rates are also not welfare-improving. This requires solving the model for 110,000 different sets of counterfactual tariffs, which is computationally feasible.
@@ 6.2 Three Country Numerical Exercises
In a three country framework, countries are able to have bilateral trade deficits even when there are no aggregate trade deficits. In our numerical exercises, we generate scenarios with varying levels of bilateral trade deficits where each trade deficit (surplus) is perfectly offset by a trade surplus (deficit) with the other trading partner. Specifically, for a bilateral deficit of $d$ between Country 1 and Country 2, we subtract $d/2$ from Country 3’s exports to Country 1 and add $d/2$ to exports from Country 1 to Country 3. Similarly, but the reverse for Country 2. This ensures that no country has an aggregate trade imbalance, despite its bilateral imbalances. We consider optimal tariff setting between Country 1 and Country 2, and assume that the Country 3 does not participate and leaves its tariff rates unchanged.
@@
When there are only bilateral deficits, and no aggregate deficits, tariffs are effectively mirrored between Country 1 and Country 2. As Country 1 exhibits a larger bilateral trade deficit with Country 1, its Nash equilibrium tariff rate increases while Country 2’s Nash equilibrium tariff rate decreases. This might seem to suggest that a country with a bilateral deficit will always have higher optimal tariffs than its trading partner; however, this need not be true in general. Optimal tariff rates interact not only with the bilateral imbalance between partners, but also bilateral and aggregate imbalances with further trading partners.
@@
Figure 5 presents the Nash equilibrium tariffs and resulting welfare changes under the scenario where Country 1 has a trade deficit with Country 3, and Country 2 has a trade surplus with Country 3. Instead of adjusting trade with Country 3 by $d/2$, we adjust it in the opposite direction by a fixed amount. [^22] For this scenario, we set the deficit between Country 1 and Country 3 to 15, and the same for the surplus between Country 2 and Country 3. Recall we set income, net of deficits, equal to 100 for these examples. Therefore, Country 1 has an aggregate trade deficit greater than its bilateral trade deficit with Country 2. Under this scenario, optimal tariffs are no longer mirrored; and in fact, Nash equilibrium tariffs for Country 2 increase as the bilateral deficit increases between Country 1 and Country 2, until they are higher than Country 1’s tariffs. This scenario more closely mirrors the economic situation between the United States and China, where the United States maintains a deficit with the rest of the world and China maintains a surplus. This helps explain why, under several model specifications, we will find that China’s Nash equilibrium tariffs are higher than those of the United States’, despite China having a trade surplus with the United States. It continues to be true, however, that a larger trade deficit increases the welfare of the country with the deficit in a trade war, and lowers the welfare of the country with the surplus.
@@
Although our focus is on optimal tariffs rates, it is important to note that trade imbalances do not just affect optimal tariff rates. Trade imbalances also alter the welfare implications of arbitrary tariffs. Figure 6 highlights this point by showing how the welfare of Country 1 changes as its trade balance varies when Country 1 places a flat 10 percent tariff on imports from Country 2. We consider both the case where Country 1 places unilateral tariff and where Country 2 retaliates by putting a flat 10 percent tariff on imports from Country 1. We find that Country 1’s welfare increases monotonically as its bilateral trade deficit grows (or bilateral trade surplus shrinks), for both unilateral and retaliatory 10 percent tariffs. [^23] We ignore the possibility of autarky in Figure 6, which we discuss in Section 6.1, and here focus on the scenario where the trade imbalances cannot be strategically avoided. We show these results hold in both the scenario with two countries, where bilateral imbalances are aggregate imbalances, and with three countries, where bilateral imbalances can exist with or without aggregate imbalances.
@@
Table 2 presents optimal unilateral and Nash equilibrium tariffs for each of the 22 tradable goods-producing sectors (unilateral tariffs are reported only for the United States). Results are reported for changes relative to two distinct baseline economies. The first scenario is when the model is calibrated to match exactly the 2014 baseline economy. This baseline was chosen because it is the most recent year for which we have input-output data available, and because it is two years prior to the election of former President Trump, who used the executive powers granted through his office to initiate the first round of tariff increases for U.S. imports of goods from China. The second baseline we consider is a world economy with free trade between all bilateral country pairs, including the United States and China. We construct the free trade baseline by calibrating the model to match 2014 data and computing the resulting counterfactual world economy after setting all tariffs equal to 1 for all country pairs and sectors. For each baseline, we solve for optimal unilateral and Nash equilibrium tariffs that result between the United States and China. Optimal tariffs are non-uniform because sectoral elasticities vary across sectors and, in the pre-trade war baseline, partner tariffs vary by sector. The table also reports the trade weighted average tariff rate, $\overline{\tau^{*}}$, for each baseline, where the weights for each sector are equal to trade values in the baseline data.
@@
Table 3 presents the optimal tariff rates and counterfactual welfare changes for a model without services or input-output linkages. The United States experiences welfare gains relative to the pre-trade war baseline in both frameworks, however, excluding services and input-output linkages makes it appear as if the United States would benefit from a trade war even relative to a free trade baseline. Compared to this simpler framework, including services and input-output linkages reduces the welfare gains of a trade war from the pre-trade war baseline by 81.8 percent for the United States (from 0.044 percent to 0.008 percent) and increases the welfare losses for China by more than double (133.8 percent; from a loss of 0.059 percent to a loss of 0.138 percent).

### v2

@@ 4 Quantitative Model
where ${t_{ij}^{s}}$ is an iceberg trade cost, for example, due to non-tariff trade barriers or freight costs, and $z_{j}^{s}$ is productivity. This means labor is given by
@@
The total supply of goods is consumption plus the intermediate inputs used in each sector,
$\sum_{s=1}^{S}\sum_{i=1}^{J}\left(x_{ij}^{sk}\right)+c_{j}^{k}=y_{j}^{k}.$ (23)
@@ 5.1 Data on Flows and Tariffs
We use the 2016 Revision of the World Input-Output Database (WIOD), as described by Timmer et al. (2015) and Timmer et al. (2016), for data on industry-level trade flows and input-output flows. This dataset includes origin-destination-sector expenditures, $p_{ij}^{s}y_{ij}^{s}/\tau_{ij}^{s}$, net of tariffs, and sector-to-sector input-output flows, $p_{j}^{k}x_{j}^{sk}$. Note that in our model, $p_{ij}^{s}y_{ij}^{s}$  is expenditures from the consumer side, which include tariffs. The flows in the WIOD data do not include tariffs, therefore we must add them to obtain tariff-inclusive expenditures. We use data on tariff rates from the United Nations’ Trade Analysis Information System (TRAINS) database. We use effectively applied weighted average tariff rates as our measure for baseline tariffs, $\tau_{ij}^{s}$. Table 1 reports pre-trade war tariff rates for each of our tradable goods-producing industries. [^17] Construction is included the Goods-Producing Industries super-sector classification by the United States’ Bureau of Labor Statistics. We exclude construction from our grouping of tradable goods-producing industries, and include it as a service industry. Service industries are still tradable, but they are not subject to tariffs. The column "USA" reports tariff rates enacted by the United States for imports from China, and the column "CHN" reports tariff rates enacted by China for imports from the United States. [^18] The tariff data provided by TRAINS is available at the 2-digit International Standard Industrial Classification of All Economic Activities (ISIC) Rev 3 level, which we convert to match the ISIC Rev. 4 sectors in the WIOD using the concordance provided by Pahl et al. (2022).
@@ 5.2 Calibration
We take estimates of the sectoral elasticities, $\sigma_{s}$, from Fontagné et al. (2022), who provide sectoral elasticity estimates for each of the tradable WIOD sectors. Their estimates are based on product-level variation in tariff rates and trade-flows, which they pool by sector to estimate sectoral elasticities. They are unable to estimate elasticities for service sectors, because trade in service sectors is not subject to tariffs. We set elasticities for service industries equal to 5.5, which is near the median value of their product-level trade elasticities.
@@
We calibrate the aggregate trade balance directly from equation (22) and initial tariff revenues from equation (21). Using equation (13), we compute income, ${I}_{i}$, as
${I}_{i}=\sum_{s=1}^{S}p_{i}^{s}c_{i}^{s}=\sum_{s=1}^{S}\sum_{i=1}^{J}p_{ij}^{s}y_{ij}^{s}-\sum_{s=1}^{S}p_{i}^{s}x_{i}^{s},$ (25)
@@ 5.3 Computation of Optimal Tariffs
The application motivating our analysis is a tariff war between two countries with a bilateral trade imbalance, therefore, we focus our analysis on optimal tariff setting between two trading partners, $i$ and $j$. To compute optimal tariffs, we solve the problem of a government choosing tariffs with a specific trading partner to maximize domestic welfare. Specifically, the government for country $i$ chooses tariffs in each sector, $\boldsymbol{\tau}_{i}\equiv\{\tau_{ij}^{s}\}_{s=1}^{S}$, to solve
@@
In general, there is no analytic solution for best response tariffs or Nash tariffs. Consequently, standard solution methods require evaluating counterfactual welfare for potential tariff candidates to find the candidates that numerically yield the highest welfare, which necessitates solving the model a large number of times. Past quantitative work has dealt with this computational intensity by focusing on small models, for example with one or two sectors (see He et al., 2017; or Chattopadhyay and Mitka, 2019), by focusing on environments where optimal tariff setting is either shown to be, or restricted to be, uniform across sectors (Caliendo et al., 2023); or by characterizing properties of optimal tariffs without calculating specific values (Costinot et al., 2020). The general approach to the problem of numerically solving for optimal tariff rates is given by Algorithm 1 (A more detailed approach is described by Algorithm C.1 in Appendix C). Running Algorithm 1 a single time with baseline tariffs yields optimal unilateral tariffs.
@@
Huang et al. (2013) show that, for certain parameterizations, it is possible to have multiple Nash equilibria in a simple two country two good CES framework. Particularly with low trade elasticities, they demonstrate the existence of Nash equilibria with extremely high tariff rates, including rates greater than 100,000 percent. For other parameterizations, the Nash equilibrium will be unique. To limit any potential impact of multiple equilibria, we restrict our analysis to tariff rates capped above at 400 percent. This rate exceeds the 99th percentile of industry level tariff rates in the TRAINS data, which is approximately 55 percent. This cap is also higher than the maximum reported ad valorem tariff rate of 350 percent set on an 8-digit Harmonized System product code by the United States in both 2014 and 2023, according to the United States International Trade Commission’s Tariff Database. [^21] https://dataweb.usitc.gov/tariff/annual We confirm that moderately lowering or raising this upper bound does not affect our quantitative results.
@@
We discuss our implementation of a hybrid genetic algorithm in Appendix C. As noted previously, there is nothing inherently unique about our chosen approach, and feasible alternatives exist, including random search, simulated annealing (Wu and Wang, 1998), particle swarm optimization (Kennedy and Eberhart, 1995), and hyperband optimization (Li et al., 2018). It should be noted that grid search is implausible due to the computational unfeasibility of evaluating all points on the grid, even for small grids. [^23] As noted in the introduction, even a discretized solution space of only 10 possible tariff rates for each sector yields 10 Sextillion possible combinations. We do not provide a detailed comparison on the relative performances of various global optimization heuristics, in part because each of them can be modified and adapted to perform better on specific problems. For instance, the genetic algorithm described in Appendix C will outperform a generic implementation of a genetic algorithm in terms of both accuracy and speed when applied to our problem of computing optimal tariff rates.
After convergence, we verify through exhaustive search that there are no single-sector deviations from Nash equilibrium tariffs that improve welfare. We evaluate counterfactual welfare changes for all tariffs between 0 and 25 percent to two decimal places for each sector. This encompasses the range of optimal tariffs we find, and we verify separately with a coarser grid that deviations with higher tariff rates are also not welfare-improving. This requires solving the model for 110,000 different sets of counterfactual tariffs, which is computationally feasible.
@@ 6.2 Three Country Numerical Exercises
In a three country framework, countries are able to have bilateral trade deficits even when there are no aggregate trade deficits. In our numerical exercises, we generate scenarios with varying levels of bilateral trade deficits where each trade deficit (surplus) is perfectly offset by a trade surplus (deficit) with the other trading partner. Specifically, for a bilateral deficit of $d$ between Country 1 and Country 2, we subtract $d/2$ from Country 3’s exports to Country 1 and add $d/2$ to exports from Country 1 to Country 3. We apply a mirrored adjustment for Country 2’s trade flows with Country 3. This ensures that no country has an aggregate trade imbalance, despite its bilateral imbalances. We consider optimal tariff setting between Country 1 and Country 2, and assume that Country 3 does not participate and leaves its tariff rates unchanged.
@@
When there are only bilateral deficits, and no aggregate deficits, tariffs are effectively mirrored between Country 1 and Country 2. As Country 1 exhibits a larger bilateral trade deficit with Country 2, its Nash equilibrium tariff rate increases while Country 2’s Nash equilibrium tariff rate decreases. This might seem to suggest that a country with a bilateral deficit will always have higher optimal tariffs than its trading partner; however, this need not be true in general. Optimal tariff rates interact not only with the bilateral imbalance between partners, but also bilateral and aggregate imbalances with further trading partners.
@@
Figure 5 presents the Nash equilibrium tariffs and resulting welfare changes under the scenario where Country 1 has a trade deficit with Country 3, and Country 2 has a trade surplus with Country 3. Instead of adjusting trade with Country 3 by $d/2$, we adjust it in the opposite direction by a fixed amount. [^24] For this scenario, we set the deficit between Country 1 and Country 3 to 15, and the same for the surplus between Country 2 and Country 3. Recall we set income, net of deficits, equal to 100 for these examples. Therefore, Country 1 has an aggregate trade deficit greater than its bilateral trade deficit with Country 2. Under this scenario, optimal tariffs are no longer mirrored; and in fact, Nash equilibrium tariffs for Country 2 increase as the bilateral deficit increases between Country 1 and Country 2, until they are higher than Country 1’s tariffs. This scenario more closely mirrors the economic situation between the United States and China, where the United States maintains a deficit with the rest of the world and China maintains a surplus. This helps explain why, under several model specifications, we will find that China’s Nash equilibrium tariffs are higher than those of the United States, despite China having a trade surplus with the United States. It continues to be true, however, that a larger trade deficit increases the welfare of the country with the deficit in a trade war, and lowers the welfare of the country with the surplus.
@@
Although our focus is on optimal tariffs rates, it is important to note that trade imbalances do not just affect optimal tariff rates. Trade imbalances also alter the welfare implications of arbitrary tariffs. Figure 6 highlights this point by showing how the welfare of Country 1 changes as its trade balance varies when Country 1 places a flat 10 percent tariff on imports from Country 2. We consider both the case where Country 1 places unilateral tariff and where Country 2 retaliates by putting a flat 10 percent tariff on imports from Country 1. We find that Country 1’s welfare increases monotonically as its bilateral trade deficit grows (or its bilateral trade surplus shrinks), for both unilateral and retaliatory 10 percent tariffs. [^25] We ignore the possibility of autarky in Figure 6, which we discuss in Section 6.1, and here focus on the scenario where the trade imbalances cannot be strategically avoided. We show these results hold in both the scenario with two countries, where bilateral imbalances are aggregate imbalances, and with three countries, where bilateral imbalances can exist with or without aggregate imbalances.
@@
Table 2 presents optimal unilateral and Nash equilibrium tariffs for each of the 22 tradable goods-producing sectors (unilateral tariffs are reported only for the United States). Results are reported for changes relative to two distinct baseline economies. The first scenario is when the model is calibrated to match exactly the 2014 baseline economy. This baseline was chosen because it is the most recent year for which we have input-output data available, and because it is two years prior to the initial election of President Trump, who used the executive powers granted through his office to initiate the first round of tariff increases for U.S. imports of goods from China. The second baseline we consider is a world economy with free trade between all bilateral country pairs, including the United States and China. We construct the free trade baseline by calibrating the model to match 2014 data and computing the resulting counterfactual world economy after setting all tariffs equal to 0 ($\tau^{s}_{ij}=1$) for all country pairs $ij$ and sectors $s$. For each baseline, we solve for optimal unilateral and Nash equilibrium tariffs between the United States and China. Optimal tariffs are non-uniform because sectoral elasticities vary across sectors and, in the pre-trade war baseline, partner tariffs vary by sector. The table also reports the trade-weighted average tariff rate, $\overline{\tau^{*}}$, for each baseline, where the weights for each sector are equal to trade values in the baseline data.